how do you find the answer to 5x-2x=54, it would be very appreciated with a long explained answer, thank you

Answer by ReadingBoosters(3246)

Starting with
5x-2x=54
Combine the x variables by subtracting the coefficients 5 and 2
5x-2x = 3x
Now, you are left with
3x=54
Divide both sides by 3 to isolate the x variable
3x/3 = 54/3
x = 18
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com